x=acost y=bsint then dsquarey by dxsquare at t=pie/4

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
x^2 / a^2 + y^2 /b^2 = 1
 
now
 
2x /a^2 + 2y dy/dx / b^2 = 0
 
- (x/y ) (b^2 /a^2) = dy/dx
 
- b^2 / a^2 = y d2y /dx2 + (dy/dx)^2
 
now  find y and dy/dx at t= pi/4